# Nonholonomic Robot Parking by Feedback— Part II: Nonmodular, Inverse Optimal, Adaptive, Prescribed/Fixed-Time and Safe Designs

### v1

@@ 1.2 Contributions and Organization
In Section 4 we exploit our construction of globally strict CLFs in [38] to design an extensive family of inverse optimal controllers, including linear, sublinear, bounded, and “relay-approximating” feedback laws, providing a systematic way to trade off control effort and convergence behavior while preserving global stability.
@@ 3.1 Unidirectional GES BAR-FLi
with $q=\sqrt{k_{1}/k_{3}}$ is a globally strict CLF for (2) on $\mathcal{S}_{2}$ with respect to the input pair $(v/\rho,\omega)$ in the sense of [38, Def. 2].
@@ 3.2 Bidirectional GES Backstepping
When the vehicle is allowed to reverse the forward velocity, it would be more efficient to exploit this capability when needed, especially in parking scenarios. The following Theorem exploits this and achieves GES. Fig. 2 compares the trajectory generated by the controller from Theorem 2 with that of [30, Sec. III.A].
@@
with $q=\sqrt{k_{1}/k_{3}}$ is a globally strict CLF for (2) on $\mathcal{S}$ with respect to the input pair $(v/\rho,\omega)$ in the sense of [38, Def. 2].
@@
Figure 2: Comparison of the control law from Theorem 2 (blue) and [30, Sec. III.A] (red) with initial conditions $(\rho_{0},\delta_{0},\gamma_{0})=(1,-4\pi/5,\pi)$, showing that restricting to unidirectional forward velocity yields a less efficient trajectory to the target.
@@ 4 Inverse Optimal Design
We now turn to the design of optimal controllers. The inescapable ‘curse of dimensionality’ of classical optimal control is to an equal degree about storing the results as it is about solving the HJB PDE. For instance, on devices with less than a terabyte of storage, there is not enough space to store the HJB solution on a uniformly quantized grid for a robotic manipulator with more than 3DOF. Instead, in this section, we derive a family of inverse optimal controllers for the unicycle based on the numerous CLFs available. It is important to note that the derived controllers fall into the category of damping $L_{g}V$ controllers (i.e., see [15], [36, Prop. 5.9.1]) that immediately satisfy the Jurdjevic-Quinn conditions and globally stabilize the unicycle.
@@ 4.1 Alternative controllers: $L_{g}V$ (gradient) designs
This property holds for the following reason. For systems affine in control, $\dot{x}=f(x)+g(x)u$, a positive definite radially unbounded differentiable function $V(x)$ is a CLF if and only if the following implication holds: for all $x\neq 0$ where $L_{g}V(x)=0$ it holds that $L_{f}V(x)<0$. When $f(x)\equiv 0$, this implication gives that $L_{g}V(x)\neq 0$ for all $x\neq 0$, namely, that $L_{g}V(x)=0$ only at $x=0$. For instance, we establish that all Lyapunov functions in [38, Thm. 1] are CLFs in the sense of [38, Def. 2] with the Genova feedback law.
While our theorems establish that we construct Lyapunov functions $V(\rho,\delta,\gamma)$ such that $|L_{\bar{g}}V(\Xi)|$ is positive definite in $\Xi$, it is not radially unbounded. Radial unboundedness can be achieved with a modification of the CLF, which is described in the final segment of the proof of [22, Thm. 3.2].
As an illustration of what the Lie derivatives $L_{\bar{g}_{1}}V$ and $L_{g_{2}}V$ look like, we compute them for the Genova CLF in [38, Eq. (25) with (13) and (24)], with unity gains:
@@
Similarly, the Lie derivatives $L_{\bar{g}_{1}}V$ and $L_{g_{2}}V$ for the GloBa CLF in [38, Eq. (37) with (13) and (38)], with unity gains, yield
@@
In both cases, the introduction of the change of control variable from $v$ to $v/\rho$ has successfully “desingularized” the Lie derivative expressions as seen in (34a) and (35a).
@@ 4.2 Basic quadratic inverse optimality
Such costs are meaningful in the sense of imposing (1) a zero penalty when all the state variables $\rho,\delta,\gamma$ and both of the ($\rho$-weighted) control inputs $v/\rho$ and $\omega$ are zero, and (2) a positive penalty when any of the states and either of the (weighted) controls are nonzero. Note that when $\rho=0$, the forward velocity input is subjected to an infinite penalty, which is desirable, as it discourages movement away from the positional origin once it is reached.
It is relevant to examine the complexity of the $L_{g}V$ controllers as compared to their counterparts used to derive the CLFs. To that end, we focus on the GloBa controller and note that the inverse optimal steering feedback (36b) with (35b) is much simpler than the non-optimal GloBa feedback [38, Eq. (7) with (41)]. But, in contrast, compared to the basic forward velocity feedback $v=\rho\cos\gamma$, the inverse optimal forward velocity feedback (36) with (35a), given by $v^{*}=\rho^{3}\cos\gamma-\rho\sin\gamma\left[\delta+\left(1+\frac{1}{1+4\delta^{2}}\right)z\right]$ is considerably more complex and depends not only on $(\rho,\gamma)$ but also on $\delta$ (when the vehicle points sidewise relative to the target, the velocity is nonzero).
@@ 4.3 General inverse optimal designs
The transformation has useful properties, given in [22, Lemma A1]. Also, to minimize notational repetition, we use the subscript $i$ in place of the subscripts $1$ and $2$, corresponding to the vector fields $\bar{g}_{1}$ and $g_{2}$ and their associated control inputs $v/\rho$ and $\omega$.
@@
with $\nu_{1}(\Xi)\coloneqq L_{\bar{g}_{1}}V$ and $\nu_{2}(\Xi)\coloneqq L_{g_{2}}V$ denoted for the CLFs given by [38, Thms. 1-8], and for all initial conditions on the respective state spaces $\mathcal{S},\mathcal{S}_{1},\mathcal{S}_{2},$ and $\mathcal{S}_{3}$. Additionally, the feedback law
@@
is continuous in $(\nu_{1},\nu_{2})$ and renders the point $\rho=\delta=\gamma=0$ of the system (40) GAS on the respective state spaces $\mathcal{S},\mathcal{S}_{1},\mathcal{S}_{2},$ and $\mathcal{S}_{3}$, in accordance to [38, Def. 1].
@@
By [22, Lemma A1], $\ell\eta_{i}$ are class $\mathcal{K}_{\infty}$ functions, and since $V$ is a strict CLF, it follows by definition that $\nu_{1}(\Xi)=0$ and $\nu_{2}(\Xi)=0$ if and only if $\Xi=0$. Hence, (47) holds, and the point $\rho=\delta=\gamma=0$ is GAS under each CLF on its respective state space when using the feedback law (45).
@@
From [22, Lemma A1], we see that $r(\eta^{\prime})^{-1}(r)=\ell\eta(r)+\eta((\eta^{\prime})^{-1}(r))$ and (48) can be rewritten as
@@ 4.4 Specific inverse optimal choices
Additionally, since inverse optimal redesign relies on $L_{g}V$-based controllers, different CLFs naturally lead to different system behaviors. Hence, we conduct simulation comparisons between the examples using a particular GloBa CLF in the following corollary derived from [38, Thm. 7]:
@@
with $k_{1},k_{2},k_{3}>0$ is a globally strict CLF on $\mathcal{S}$ for (2) with respect to the input pair $(v/\rho,\omega)$ in the sense of [38, Def. 2].
@@
Figure 7: Parallel parking trajectories for forward $L_{g}V$ control (red), reverse $L_{g}V$ control (blue), forward GloBa control [38, Thm. 7] (cyan), and reverse BoLSA control [38, Thm. 2] (yellow) with opacity representing different gain choices.
@@
Parallel parking has long been considered a benchmark maneuver for nonholonomic vehicles due to its intrinsic difficulty requiring precise coordination of orientation and position, often in confined spaces. It captures the core difficulty of planning and control for systems that cannot move freely in all directions, such as cars or mobile robots. With a wide range of control laws and infinitely many configurable parameters in both this paper and [38], our framework offers numerous options that can be tailored to the specific needs of a given task. Fig 7 highlights several representative examples from this toolbox. The $L_{g}V$ controllers exhibit behaviors that closely resemble human-like parallel parking, with the reverse $L_{g}V$ controller exactly representing the traditional approach used by drivers. In contrast, the standard GloBa and reversed BoLSA controllers [38, Thms. 2 and 7] produce smoother “S”-shaped trajectories that are more appropriate for maneuvering in tighter spaces.
@@ 5 Adaptive Stabilization under Unknown Input Coefficients
Consider the system (73) with arbitrary unknown $b_{1},b_{2}>0$, along with the control law (75), (76) and the update laws (80), (81). For each CLF $V$ given by [38, Thms. 1-8], with class $\mathcal{K}_{\infty}$ functions $\alpha_{1},\alpha_{2}$ such that
@@
To illustrate the performance of the proposed method, we consider the BoFo CLF $V=\rho^{2}+V_{\delta\gamma}$, where $V_{\delta\gamma}$ is defined in [38, Thm. 6], and choose gains $k_{1}=3$, $k_{2}=3.5$, and $k_{3}=4$. In Fig. 8, we compare the quadratic cost $L_{g}V$ controller (red), given by $v=-\rho L_{\bar{g}_{1}}V$ and $\omega=-L_{g_{2}}V$, with the adaptive control law in (80)–(81), which uses a normalization function of the form $n(V)=V$, initial parameter estimates $\hat{\varepsilon}_{1}(0)=\hat{\varepsilon}_{2}(0)=0$, and adaptation gains $\mu_{1}=\mu_{2}=0.5$ (blue) and $\mu_{1}=\mu_{2}=1$ (cyan). The system is initialized at $(\rho(0),\delta(0),\gamma(0))=(1,-\pi/2,-\pi/2)$, under uncertainty in $b_{1}=b_{2}=1$. For the quadratic cost $L_{g}V$ controller, this corresponds to the assumption of $b_{1}\varepsilon_{1}=b_{2}\varepsilon_{2}=1$. The adaptive controller exhibited smaller peak values in both forward velocity and steering input (Figure 8(b)), even while achieving a faster decay of the state norm (Figure 8(a)). This behavior is clearly due to $\hat{\varepsilon}_{1,2}$ starting small but quickly growing beyond $1$, thereby accelerating the decay rate to be faster than the $L_{g}V$ controller.
@@ 7 Nonovershooting Control: Strictly Forward Parking without Curb Violation
We utilize the BAR-FLi CLFs as in [38, Thm. 8], which have barrier-like properties as discussed in [38, Section III.D] and serve to restrict the polar angle $\delta$. Building on the backstepping transformation of BAR-FLi, we design $\tilde{\omega}(t)$ to ensure that $\delta(t)\in[0,\pi)$ for all $t\geq 0$, which implies that $y(t)\leq 0$ for all $t\geq 0$. The following result introduces a control law that renders the origin GAS on $\mathcal{S}_{2}$, and further ensures $\delta(t)\in[0,\pi)$ for all $t\geq 0$ if the initial condition satisfies $\delta_{0}\in(0,\pi)$ and $\gamma_{0}\in(-\pi/4,\pi/4)$ through an appropriate gain selection. Furthermore, under the same initial condition constraints, the forward velocity input $v(t)$ remains nonnegative for all $t\geq 0$, which is crucial in applications where the unicycle model represents fixed-wing aircraft or guided missiles. Fig 10 illustrates the corresponding admissible region in the $(\delta,\gamma)$ space and its geometric interpretation in Cartesian coordinates.
@@
and with $k_{1},k_{2},k_{3},k_{4}>0$, render the point $\rho=\delta=\gamma=0$ GAS on $\mathcal{S}_{2}$, in accordance with [38, Def. 1]. Moreover, if the initial conditions satisfy $\delta_{0}\in(0,\pi)$ and $\gamma_{0}\in(-\pi/4,\pi/4)$ with $k_{2}$ chosen as
@@
Consider again the BAR-FLi Lyapunov function introduced in [38, Thm. 8],
@@
is negative definite on $\{\rho\geq 0\}\times\mathcal{T}_{2}$. By reasoning similar to that of the proof of [38, Thm. 8], the point $\rho=\delta=\gamma=0$ is GAS on $\mathcal{S}_{2}$, implying that $(x(t),y(t),\theta(t))\to 0$ as $t\to\infty$.

### v2

@@ 1.2 Contributions and Organization
In Section 4 we exploit our construction of globally strict CLFs in [39] to design an extensive family of inverse optimal controllers, including linear, sublinear, bounded, and “relay-approximating” feedback laws, providing a systematic way to trade off control effort and convergence behavior while preserving global stability. Portions of the results presented without proofs appear in [18].
@@ 3.1 Unidirectional GES BAR-FLi
with $q=\sqrt{k_{1}/k_{3}}$ is a globally strict CLF for (2) on $\mathcal{S}_{2}$ with respect to the input pair $(v/\rho,\omega)$ in the sense of [39, Def. 2].
@@ 3.2 Bidirectional GES Backstepping
When the vehicle is allowed to reverse the forward velocity, it would be more efficient to exploit this capability when needed, especially in parking scenarios. The following Theorem exploits this and achieves GES. Fig. 2 compares the trajectory generated by the controller from Theorem 2 with that of [31, Sec. III.A].
@@
with $q=\sqrt{k_{1}/k_{3}}$ is a globally strict CLF for (2) on $\mathcal{S}$ with respect to the input pair $(v/\rho,\omega)$ in the sense of [39, Def. 2].
@@
Figure 2: Comparison of the control law from Theorem 2 (blue) and [31, Sec. III.A] (red) with initial conditions $(\rho_{0},\delta_{0},\gamma_{0})=(1,-4\pi/5,\pi)$, showing that restricting to unidirectional forward velocity yields a less efficient trajectory to the target.
@@ 4 Inverse Optimal Design
We now turn to the design of optimal controllers. The inescapable ‘curse of dimensionality’ of classical optimal control is to an equal degree about storing the results as it is about solving the HJB PDE. For instance, on devices with less than a terabyte of storage, there is not enough space to store the HJB solution on a uniformly quantized grid for a robotic manipulator with more than 3DOF. Instead, in this section, we derive a family of inverse optimal controllers for the unicycle based on the numerous CLFs available. It is important to note that the derived controllers fall into the category of damping $L_{g}V$ controllers (i.e., see [15], [37, Prop. 5.9.1]) that immediately satisfy the Jurdjevic-Quinn conditions and globally stabilize the unicycle.
@@ 4.1 Alternative controllers: $L_{g}V$ (gradient) designs
This property holds for the following reason. For systems affine in control, $\dot{x}=f(x)+g(x)u$, a positive definite radially unbounded differentiable function $V(x)$ is a CLF if and only if the following implication holds: for all $x\neq 0$ where $L_{g}V(x)=0$ it holds that $L_{f}V(x)<0$. When $f(x)\equiv 0$, this implication gives that $L_{g}V(x)\neq 0$ for all $x\neq 0$, namely, that $L_{g}V(x)=0$ only at $x=0$. For instance, we establish that all Lyapunov functions in [39, Thm. 1] are CLFs in the sense of [39, Def. 2] with the Genova feedback law.
While our theorems establish that we construct Lyapunov functions $V(\rho,\delta,\gamma)$ such that $|L_{\bar{g}}V(\Xi)|$ is positive definite in $\Xi$, it is not radially unbounded. Radial unboundedness can be achieved with a modification of the CLF, which is described in the final segment of the proof of [23, Thm. 3.2].
As an illustration of what the Lie derivatives $L_{\bar{g}_{1}}V$ and $L_{g_{2}}V$ look like, we compute them for the Genova CLF in [39, Eq. (25) with (13) and (24)], with unity gains:
@@
Similarly, the Lie derivatives $L_{\bar{g}_{1}}V$ and $L_{g_{2}}V$ for the GloBa CLF in [39, Eq. (37) with (13) and (38)], with unity gains, yield
@@
where $z=\gamma+\frac{1}{2}\arctan(2\delta)$. In both cases, the introduction of the change of control variable from $v$ to $v/\rho$ has successfully “desingularized” the Lie derivative expressions as seen in (34a) and (35a).
@@ 4.2 Basic quadratic inverse optimality
Such costs are meaningful in the sense of imposing (i.) a zero penalty when all the state variables $\rho,\delta,\gamma$ and both of the ($\rho$-weighted) control inputs $v/\rho$ and $\omega$ are zero, and (ii.) a positive penalty when any of the states and either of the (weighted) controls are nonzero. Note that when $\rho=0$, the forward velocity input is subjected to an infinite penalty, which is desirable, as it discourages movement away from the positional origin once it is reached.
It is relevant to examine the complexity of the $L_{g}V$ controllers as compared to their counterparts used to derive the CLFs. To that end, we focus on the GloBa controller and note that the inverse optimal steering feedback (36b) with (35b) is much simpler than the non-optimal GloBa feedback [39, Eq. (7) with (41)]. But, in contrast, compared to the basic forward velocity feedback $v=\rho\cos\gamma$, the inverse optimal forward velocity feedback (36) with (35a), given by $v^{*}=\rho^{3}\cos\gamma-\rho\sin\gamma\left[\delta+\left(1+\frac{1}{1+4\delta^{2}}\right)z\right]$ is considerably more complex and depends not only on $(\rho,\gamma)$ but also on $\delta$ (when the vehicle points sidewise relative to the target, the velocity is nonzero).
@@ 4.3 General inverse optimal designs
The transformation has useful properties, given in [23, Lemma A1]. Also, to minimize notational repetition, we use the subscript $i$ in place of the subscripts $1$ and $2$, corresponding to the vector fields $\bar{g}_{1}$ and $g_{2}$ and their associated control inputs $v/\rho$ and $\omega$.
@@
with $\nu_{1}(\Xi)\coloneqq L_{\bar{g}_{1}}V$ and $\nu_{2}(\Xi)\coloneqq L_{g_{2}}V$ denoted for the CLFs given by [39, Thms. 1-8], and for all initial conditions on the respective state spaces $\mathcal{S},\mathcal{S}_{1},\mathcal{S}_{2},$ and $\mathcal{S}_{3}$. Additionally, the feedback law
@@
is continuous in $(\nu_{1},\nu_{2})$ and renders the point $\rho=\delta=\gamma=0$ of the system (40) GAS on the respective state spaces $\mathcal{S},\mathcal{S}_{1},\mathcal{S}_{2},$ and $\mathcal{S}_{3}$, in accordance to [39, Def. 1].
@@
By [23, Lemma A1], $\ell\eta_{i}$ are class $\mathcal{K}_{\infty}$ functions, and since $V$ is a strict CLF, it follows by definition that $\nu_{1}(\Xi)=0$ and $\nu_{2}(\Xi)=0$ if and only if $\Xi=0$. Hence, (47) holds, and the point $\rho=\delta=\gamma=0$ is GAS under each CLF on its respective state space when using the feedback law (45).
@@
From [23, Lemma A1], we see that $r(\eta^{\prime})^{-1}(r)=\ell\eta(r)+\eta((\eta^{\prime})^{-1}(r))$ and (48) can be rewritten as
@@ 4.4 Specific inverse optimal choices
Additionally, since inverse optimal redesign relies on $L_{g}V$-based controllers, different CLFs naturally lead to different system behaviors. Hence, we conduct simulation comparisons between the examples using a particular GloBa CLF in the following corollary derived from [39, Thm. 7]:
@@
with $k_{1},k_{2},k_{3}>0$ is a globally strict CLF on $\mathcal{S}$ for (2) with respect to the input pair $(v/\rho,\omega)$ in the sense of [39, Def. 2].
@@
Figure 7: Parallel parking trajectories for forward $L_{g}V$ control (red), reverse $L_{g}V$ control (blue), forward GloBa control [39, Thm. 7] (cyan), and reverse BoLSA control [39, Thm. 2] (yellow) with opacity representing different gain choices.
@@
Parallel parking has long been considered a benchmark maneuver for nonholonomic vehicles due to its intrinsic difficulty requiring precise coordination of orientation and position, often in confined spaces. It captures the core difficulty of planning and control for systems that cannot move freely in all directions, such as cars or mobile robots. With a wide range of control laws and infinitely many configurable parameters in both this paper and [39], our framework offers numerous options that can be tailored to the specific needs of a given task. Fig 7 highlights several representative examples from this toolbox. The $L_{g}V$ controllers exhibit behaviors that closely resemble human-like parallel parking, with the reverse $L_{g}V$ controller exactly representing the traditional approach used by drivers. In contrast, the standard GloBa and reversed BoLSA controllers [39, Thms. 2 and 7] produce smoother “S”-shaped trajectories that are more appropriate for maneuvering in tighter spaces.
@@ 5 Adaptive Stabilization under Unknown Input Coefficients
Consider the system (73) with arbitrary unknown $b_{1},b_{2}>0$, along with the control law (75), (76) and the update laws (80), (81). For each CLF $V$ given by [39, Thms. 1-8], with class $\mathcal{K}_{\infty}$ functions $\alpha_{1},\alpha_{2}$ such that
@@
To illustrate the performance of the proposed method, we consider the BoFo CLF $V=\rho^{2}+V_{\delta\gamma}$, where $V_{\delta\gamma}$ is defined in [39, Thm. 6], and choose gains $k_{1}=3$, $k_{2}=3.5$, and $k_{3}=4$. In Fig. 8, we compare the quadratic cost $L_{g}V$ controller (red), given by $v=-\rho L_{\bar{g}_{1}}V$ and $\omega=-L_{g_{2}}V$, with the adaptive control law in (80)–(81), which uses a normalization function of the form $n(V)=V$, initial parameter estimates $\hat{\varepsilon}_{1}(0)=\hat{\varepsilon}_{2}(0)=0$, and adaptation gains $\mu_{1}=\mu_{2}=0.5$ (blue) and $\mu_{1}=\mu_{2}=1$ (cyan). The system is initialized at $(\rho(0),\delta(0),\gamma(0))=(1,-\pi/2,-\pi/2)$, under uncertainty in $b_{1}=b_{2}=1$. For the quadratic cost $L_{g}V$ controller, this corresponds to the assumption of $b_{1}\varepsilon_{1}=b_{2}\varepsilon_{2}=1$. The adaptive controller exhibited smaller peak values in both forward velocity and steering input (Figure 8(b)), even while achieving a faster decay of the state norm (Figure 8(a)). This behavior is clearly due to $\hat{\varepsilon}_{1,2}$ starting small but quickly growing beyond $1$, thereby accelerating the decay rate to be faster than the $L_{g}V$ controller.
@@ 7 Nonovershooting Control: Strictly Forward Parking without Curb Violation
We utilize the BAR-FLi CLFs as in [39, Thm. 8], which have barrier-like properties as discussed in [39, Section III.D] and serve to restrict the polar angle $\delta$. Building on the backstepping transformation of BAR-FLi, we design $\tilde{\omega}(t)$ to ensure that $\delta(t)\in[0,\pi)$ for all $t\geq 0$, which implies that $y(t)\leq 0$ for all $t\geq 0$. The following result introduces a control law that renders the origin GAS on $\mathcal{S}_{2}$, and further ensures $\delta(t)\in[0,\pi)$ for all $t\geq 0$ if the initial condition satisfies $\delta_{0}\in(0,\pi)$ and $\gamma_{0}\in(-\pi/4,\pi/4)$ through an appropriate gain selection. Furthermore, under the same initial condition constraints, the forward velocity input $v(t)$ remains nonnegative for all $t\geq 0$, which is crucial in applications where the unicycle model represents fixed-wing aircraft or guided missiles. Fig 10 illustrates the corresponding admissible region in the $(\delta,\gamma)$ space and its geometric interpretation in Cartesian coordinates.
@@
and with $k_{1},k_{2},k_{3},k_{4}>0$, render the point $\rho=\delta=\gamma=0$ GAS on $\mathcal{S}_{2}$, in accordance with [39, Def. 1]. Moreover, if the initial conditions satisfy $\delta_{0}\in(0,\pi)$ and $\gamma_{0}\in(-\pi/4,\pi/4)$ with $k_{2}$ chosen as
@@
Consider again the BAR-FLi Lyapunov function introduced in [39, Thm. 8],
@@
is negative definite on $\{\rho\geq 0\}\times\mathcal{T}_{2}$. By reasoning similar to that of the proof of [39, Thm. 8], the point $\rho=\delta=\gamma=0$ is GAS on $\mathcal{S}_{2}$, implying that $(x(t),y(t),\theta(t))\to 0$ as $t\to\infty$.